OpenStudy (anonymous):

(1/7)(z+2)-(1/2)(z+3)=0 solve

OpenStudy (saifoo.khan):

\[(\frac17)(z+2)-\frac12(z+3) = 0\]

OpenStudy (anonymous):

2z + 4 - 7z - 21 = -5z -17

OpenStudy (anonymous):

\[ \left\{\left\{z= -\frac{17}{5}\right\}\right\} \]